OpenStudy (anonymous):

identify the slop (y=3x+4)

OpenStudy (anonymous):

slope*

OpenStudy (anonymous):

y=mx+b the slop:3

OpenStudy (anonymous):

what is the slope of (6,13) and (13,6) using y=mx+b

OpenStudy (anonymous):

\[m=\frac{ y _{2}-y _{1} }{ x _{2}-x _{1} }\] [\[m=\frac{ 6-13 }{ 13-6 }=\frac{ -7 }{ 7 }=-1\] y-y1=m(x-x1) y-13=-1(x-6) y-13=-x+6 y=-x+19